OpenStudy (jakyfraze):

What is the value of 324? A. 648 B. 128 C. 1,048,576 D. 18

OpenStudy (jakyfraze):

@Nnesha is this the same one like the last one? i choose C but I'm not sure :/

Nnesha (nnesha):

is it 32 to the 4th power ?

OpenStudy (jakyfraze):

it"s something like that

Nnesha (nnesha):

oh so use `^` for exponent 32^4 so how did you get c ? 32 to the 4th power is sameas what ?

OpenStudy (jakyfraze):

like 8*8*8

Nnesha (nnesha):

yes just like what but now the power is 4

OpenStudy (jakyfraze):

so it's like 32*32*32*32?

Nnesha (nnesha):

yes right!!

OpenStudy (jakyfraze):

so the answer is C right?

Nnesha (nnesha):

yep

OpenStudy (jakyfraze):

yeahh haha. thank you

OpenStudy (jakyfraze):

so much man, thank you

OpenStudy (jakyfraze):

Solve the following: 32 ÷ 4 + 4 × 8 = ? A. 96 B. 32 C. 40 D. 42

OpenStudy (jakyfraze):

i answered A

Nnesha (nnesha):

how did you get A ?

OpenStudy (jakyfraze):

divided 32 4=8 + 4= 12*8= 96

Nnesha (nnesha):

familiar with the pemdas rules ?

OpenStudy (jakyfraze):

yeah i kinda do

Nnesha (nnesha):

PEMDAS Parentheses Exponents MD (multiply or divide ) then AS (addition and subtraction )

Nnesha (nnesha):

32/4 is 8 right so 8+ 4 × 8 = ? no what would be the next step ?

OpenStudy (jakyfraze):

40?

OpenStudy (jakyfraze):

wait hold on let me try again sorry

OpenStudy (jakyfraze):

yeah 40

Nnesha (nnesha):

yes

OpenStudy (jakyfraze):

what did i do wrong in the first place :?

Nnesha (nnesha):

you added 12+4 and then multiplied by 8 but according to the PEMDAS rules first deal with multiplication/division and then subtraction/addition first one 8 +4 times 8 = 12 times 8( wrong) 8+4 times 8 = 8+(4 times 8) = 8+32 right

OpenStudy (jakyfraze):

ohh I'm sorry i think i get it :( pemdas rules i have to follow it in order to get the right answer

Nnesha (nnesha):

yes right

OpenStudy (jakyfraze):

ohh im so sorry

OpenStudy (jakyfraze):

man, thank you so much

OpenStudy (jakyfraze):

I'm going to close this one and open a new one, so i can give you another medal

Nnesha (nnesha):

i have to go now good luck!

OpenStudy (jakyfraze):

okay, thank you

Nnesha (nnesha):

yw